OpenStudy (anonymous):

Marna wants to store one hundred fifty cubic inches of homemade pasta sauce in a food storage container. Should she choose the cylinder with a radius of 3 inches and a height of 6 inches, or the cube with a side length of 5 inches? Explain.

OpenStudy (anonymous):

@Barrelracer011

OpenStudy (anonymous):

@Destinymasha

OpenStudy (anonymous):

@bookworm00981

OpenStudy (anonymous):

@greenlegodude57

OpenStudy (anonymous):

@mule1208

OpenStudy (anonymous):

weve been thru this mule dosnt do math

OpenStudy (barrelracer011):

Cylinder - V = πr^2h = π(3 in)^2(5 in) ≈ 141.372 in^3 Cube - V = e^3 = (6 in)^2 = 216 in^3 The cylinder is too small and the cube is too big, but better to be too big than too small. She should choose the cube. Question 5: V = πr^2h 300 km^3 = 3.14(7 km)^2 h 1.949824516... = h The height is approximately 1.9 km.